OpenStudy (renzonacino):

How do you get rid of the fraction? 36/5 pi = (9)(x)

OpenStudy (agent0smith):

Why do you need to get rid of it?

OpenStudy (renzonacino):

The fraction on the left

OpenStudy (renzonacino):

My bad xD

OpenStudy (renzonacino):

To find the height of cylinder

OpenStudy (agent0smith):

That doesn't change my question

OpenStudy (renzonacino):

I forgot how to get rid of it

OpenStudy (renzonacino):

Do I need to multiply both side by 5? to get rid of the 5?

OpenStudy (agent0smith):

Are you trying to solve for x? 36/5 pi = (9)(x) You'd divide both sides by 9.

OpenStudy (renzonacino):

Yes

OpenStudy (renzonacino):

Let me check real quick

OpenStudy (renzonacino):

How do you divide it by 9?

OpenStudy (agent0smith):

Is this actually your equation, cos if not you need to clarify \[\Large \frac{ 36 \pi }{5 } = 9x\]

OpenStudy (renzonacino):

Yes, that's my actually equation

OpenStudy (agent0smith):

So divide both sides by 9. Dividing both sides by 9 is the same as multiplying by 1/9, so do that.

OpenStudy (renzonacino):

I get \[4\pi/5\]

OpenStudy (renzonacino):

I divide 36 by 9

OpenStudy (agent0smith):

yes

OpenStudy (renzonacino):

then x is 4pi/5

OpenStudy (renzonacino):

How am I going to put that to solve cylinder height? Volume=B(h)

OpenStudy (agent0smith):

I think you should start a new question, and post your actual question. Like the whole thing.

OpenStudy (renzonacino):

Ok